OpenStudy (cold3693):

Find the real of imaginary solutions of each equation by factoring. The equation is x^4 - 5x^2 + 4 = 0

OpenStudy (zehanz):

You can factor it like this: (x² + a)(x² + b). To find the a and b, remember that \(a\cdot b=4\) and \(a+b=-5\). First, try to make a list of possible a's and b's with a product of 4. There aren't that many possibilities. Then, pick the pair that gives a sum of -5 and you're done!

OpenStudy (zehanz):

Aww, all kinds of weird diamond shapes in my reply... I meant to type this: \((x^2+a)(x^2+b)=0\)

OpenStudy (zehanz):

If you've found the a and b numbers, there is another step, so there will be 4 factors in total...

OpenStudy (chosenmatt):

yup

OpenStudy (zehanz):

I guess you've found all 4 solutions!